# Neural Cellular Automata for ARC-AGI

### v1

@@ Abstract
Cellular automata and their differentiable counterparts, Neural Cellular Automata (NCA), are highly expressive and capable of surprisingly complex behaviors. This paper explores how NCAs perform when applied to tasks requiring precise transformations and few-shot generalization, using the Abstraction and Reasoning Corpus for Artificial General Intelligence (ARC-AGI) as a domain that challenges their capabilities in ways not previously explored. Specifically, this paper uses gradient-based training to learn iterative update rules that transform input grids into their outputs from the training examples and apply them to the test inputs. Results suggest that gradient-trained NCA models are a promising and efficient approach to a range of abstract grid-based tasks from ARC. Along with discussing the impacts of various design modifications and training constraints, this work examines the behavior and properties of NCAs applied to ARC to give insights for broader applications of self-organizing systems.
@@ Introduction
Designed to be challenging for AI, the strongest early approaches to the ARC tasks were based on program search, typically using a Domain-Specific Language (DSL) constructed specifically for ARC. However, this is an exponentially large search space, so deep-learning-guided search is the popular area of focus. The most successful approaches from the 2024 ARC-Prize competition were based on large language models pioneering test-time adaptation, with little progress prior to this. While plenty of other interesting directions exist, there is less focus on them—particularly on approaches that emphasize computational efficiency or do not make use of extensive pre-training of ARC-specific data. The Neural Cellular Automata approach proposed in this paper is an example of moving away from the conventional, centralized computing paradigm of deep learning.
@@
One other work, CAX, experimented with training NCAs on 1D-ARC (Faldor and Cully,, 2024; Xu et al.,, 2024), and outperformed GPT-4 on it. 1D-ARC is an adaption of the ARC benchmark to simplified one-dimensional tasks rather than 2D grids. While it does give some insight on what NCAs struggle to learn, such as counting tasks, these tasks are significantly less complex and do not represent the combination of multiple complex priors the way ARC tasks do. The original ARC benchmark is therefore used as the domain of interest in this paper.
@@
The goal of this work is not to propose a finalized architecture for ARC, but to study how a self-organizing system like an NCA behaves when tasked with ARCs abstract grid transformations. What follows is a natural implementation of NCAs adapted to ARC formats, with minimal hand-engineering. Many variations in architecture and task setup are possible, but the focus is on a straightforward implementation that serves as a natural baseline. This approach makes it possible to examine what kinds of structure, generalization, and failure arise from the model itself, without relying on external priors or domain-specific tuning. The approach draws from previous NCA work where applicable, making it easier to compare observations across domains and to carry insights in two directions: adapting known techniques to ARC, and identifying properties that may inform future NCA applications more broadly. Where relevant, variants that were tested but were found to underperform or destabilize training are noted.
@@
It is rather surprising that the model is still able to learn under highly asynchronous conditions, and in practice many solved tasks do not seem to require stochastic updates, but overall accuracy and consistency are improved across the benchmark. Thus, stochastic cell updates act as a form of regularization by filtering ’weak’ solutions that overfit to the training examples without learning anything resembling the correct, generalizable transformation. This effect also suggests that other architectural constraints that similarly help identify stronger solutions, possibly with domain-specific biases, may be promising future improvements. A related idea from NCAs for morphogenesis (Mordvintsev et al.,, 2020) is to keep track of whether each cell is alive or not: only a single seed cell starts as a alive, and dead cells are not updated. This mechanism was not found to provide any improvement in preliminary experiments and was not used in the experiments below.
@@
First, for simpler tasks that do not rely on long distance organization, such as those that are solvable within one time-step, taking loss only on the final time step does not encourage the NCA to converge and stabilize as early as possible. As a result, the intermediate states often become corrupted and have to recover, which makes generalization harder. This kind of failure is observed in as little as 15 time steps, significantly fewer than the number of steps used for morphogenesis. Meanwhile, other tasks may benefit from or require a higher number of time steps, and thus this issue is not acceptable for an approach that is generally applicable.
Second, when training on a higher number of time steps, measuring loss only on the final one can lead to a problem with gradient flow. One solution to this is to task the network to perform incremental, additive updates, as in Mordvintsev et al., (2020), rather than requiring it to output a new state at every step. However, this modification still turns out to underperform the current approach in stability and generalization with no clear benefit.
Third, small errors tend to compound over time; constant supervision mitigates this problem. Hidden channels are also still allowed to change freely, preventing the first 10 channels from having to simultaneously handle both information propagation and loss minimization, which could otherwise lead to unstable or entangled behavior. However, such training still makes convergence greedy, and the approach could miss solutions that require more complex dynamics.
@@
The models were trained for 800 epochs using the AdamW optimizer, with a learning rate linearly decayed from 0.002 to 0.0001. They have just over 10,000 parameters. In practice, much fewer than 20 hidden channels were needed for a majority of tasks, allowing models smaller than 2500 parameters to perform similarly in some cases.
@@
Indeed, the aim is not to compete with such systems, but to study a fundamentally different approach to problem-solving through local, self-organizing computation. Therefore, a closer examination of tasks will be more insightful than overall performance, as will be illustrated in the next subsection. Finally, on 95 of the attempted tasks, training achieves a categorical cross entropy loss $<0.01$, with only half of the successful tasks in this group. An additional 48 tasks had over 90% of pixels in the end grid correct. While these metrics are not perfect signals of success, they further underscore the value of deeper analysis into the behaviors and limitations of this approach.
@@
Synthesizing a program to output a spiral given these examples would not be trivial, yet this task is surprisingly simple for cellular automata. The system learned a compact, iterative rule that is precise enough to have an effective degree of size invariance, capable of scaling up to a 100$\times$100 grid (Figure 3). Encoding the solution as local state transitions demonstrates how complexity of a transformation in a NCA differs fundamentally from that in centralized models. A possible explanation for the success on this task is that it is one with a consistent initial state and relatively few transition states that need to be learned. The rules are applied iteratively and there is little structural variation or noise that lead to overfitting. The only difference between inputs is grid size, which mostly only influences how many transition steps are needed before the full grid is settled. It is not always obvious what structures or operations are more naturally and efficiently represented by the properties of a particular self-organizing system, making them an interesting direction for benchmarks like ARC and broader computational modeling.
Figure 3: Spiral pattern generalized to 100$\times$100 grid, shown at three stages of the update process (t=0, 50, 110). The model was trained only on the small examples from Figure 2 and tested with asynchronous updates on, yet maintains structural coherence across space and time, indicating strong generalization and stability.
@@
As stated previously, many of the successfully solved tasks do not rely on asynchronous updates. However, some tasks demonstrate that the regularization provided by asynchrony can make a difference.
Figure 4: Task ID: 3aa6fb7a. Input-output pairs. The left two are training pairs, and the right pair is the test task. This task is structurally simple, filling in the corner of 2$\times$2 squares, but one of four orientations does not appear in the training examples, which is a problem for Neural Cellular Automata.
@@
The choice of number of time steps is another factor with noticeable impact. For benchmarking, 10 steps were used for consistency, though this setting is somewhat arbitrary and not optimal for every task. Some tasks, like the spiral, generalized more reliably to larger grids (e.g., 100$\times$100) when trained with 30 steps, despite identical training examples. This suggests that certain transformations benefit from longer iterative computation, and that fixed step counts may limit what the model can express.
@@
A key strength of this approach is its computational efficiency. Models trained in these experiments required only minutes per task on a single consumer GPU. This contrasts sharply with ARC approaches using large language models or other deep learning architectures, which typically involve extensive pre-training or many hours of computation.
However, direct comparison with these approaches is not straightforward or particularly meaningful due to fundamental differences in assumptions and constraints. Instead of seeing these approaches as competitors, it may be more productive to consider how their insights and techniques might integrate with NCAs and potentially address limitations of the current approach.
@@
  • Architectural improvements: Exploring ideas from mainstream neural network research, such as attention, gating, or adaptive computation, that address similar weaknesses in recurrent models but have not yet been widely explored with self-organizating models. Given their structural similarities to other neural architectures, NCAs could benefit from adapting these techniques in ways that preserve local update dynamics.

### v2

@@ Abstract
Cellular automata and their differentiable counterparts, Neural Cellular Automata (NCA), are highly expressive and capable of surprisingly complex behaviors. This paper explores how NCAs perform when applied to tasks requiring precise transformations and few-shot generalization, using the Abstraction and Reasoning Corpus for Artificial General Intelligence (ARC-AGI) as a domain that challenges their capabilities in ways not previously explored. Specifically, this paper uses gradient-based training to learn iterative update rules that transform input grids into their outputs from the training examples and then applies them to the test inputs. Results suggest that gradient-trained NCA models are a promising and efficient approach to a range of abstract grid-based tasks from ARC. Along with discussing the impacts of various design modifications and training constraints, this work examines the behavior and properties of NCAs applied to ARC to give insights for broader applications of self-organizing systems.
@@ Introduction
Designed to be challenging for AI, the strongest early approaches to the ARC tasks were based on program search, typically using a Domain-Specific Language (DSL) constructed specifically for ARC. However, this is an exponentially large search space, so deep-learning-guided search is the popular area of focus. The most successful approaches from the 2024 ARC-Prize competition were based on large language models pioneering test-time adaptation, with little progress prior to this. While plenty of other interesting directions exist, there is less focus on them—particularly on approaches that emphasize computational efficiency or do not make use of extensive pretraining of ARC-specific data. The Neural Cellular Automata approach proposed in this paper is an example of moving away from the conventional, centralized computing paradigm of deep learning.
@@
One other work, CAX, experimented with training NCAs on 1D-ARC (Faldor and Cully,, 2024; Xu et al.,, 2024), and outperformed GPT-4 on it. 1D-ARC is an adaptation of the ARC benchmark to simplified one-dimensional tasks rather than 2D grids. While it does give some insight on what NCAs struggle to learn, such as counting tasks, these tasks are significantly less complex and do not represent the combination of multiple complex priors the way ARC tasks do. The original ARC benchmark is therefore used as the domain of interest in this paper.
Concurrent to this work, Guichard et al., (2025) independently applied Neural Cellular Automata to ARC, introducing an augmented architecture with hidden memory states. Their focus was on leveraging this extension to enhance the performance of standard NCAs, whereas this work seeks to understand the capabilities and limits of NCAs by examining their characteristics on ARC tasks, and establish directions for future work in self-organizing computation.
@@
The goal of this work is not to propose a finalized architecture for ARC, but to study how a self-organizing system like an NCA behaves when tasked with ARC’s abstract grid transformations. What follows is a natural implementation of NCAs adapted to ARC formats, with minimal hand-engineering. Many variations in architecture and task setup are possible, but the focus is on a straightforward implementation that serves as a natural baseline. This approach makes it possible to examine what kinds of structure, generalization, and failure arise from the model itself, without relying on external priors or domain-specific tuning. The approach draws from previous NCA work where applicable, making it easier to compare observations across domains and to carry insights in two directions: adapting known techniques to ARC, and identifying properties that may inform future NCA applications more broadly. Where relevant, variants that were tested but were found to underperform or destabilize training are noted.
@@
It is rather surprising that the model is still able to learn under highly asynchronous conditions, and in practice many solved tasks do not seem to require stochastic updates, but overall accuracy and consistency are improved across the benchmark. Thus, stochastic cell updates act as a form of regularization by filtering ’weak’ solutions that overfit to the training examples without learning anything resembling the correct, generalizable transformation. This effect also suggests that other architectural constraints that similarly help identify stronger solutions, possibly with domain-specific biases, may be promising future improvements. A related idea from NCAs for morphogenesis (Mordvintsev et al.,, 2020) is to keep track of whether each cell is alive or not: only a single seed cell starts as alive, and dead cells are not updated. This mechanism was not found to provide any improvement in preliminary experiments and was not used in the experiments below.
@@
  1. 1.
    For simpler tasks that do not rely on long distance organization, such as those that are solvable within one time-step, taking loss only on the final time step does not encourage the NCA to converge and stabilize as early as possible. As a result, the intermediate states often become corrupted and have to recover, which makes generalization harder. This kind of failure is observed in as little as 15 time steps, significantly fewer than the number of steps used for morphogenesis. Meanwhile, other tasks may benefit from or require a higher number of time steps, and thus this issue is not acceptable for an approach that is generally applicable.
  2. 2.
    When training on a higher number of time steps, measuring loss only on the final one can lead to a problem with gradient flow. One solution to this is to task the network to perform incremental, additive updates, as in Mordvintsev et al., (2020), rather than requiring it to output a new state at every step. However, this modification still turns out to underperform the current approach in stability and generalization with no clear benefit.
  3. 3.
    Small errors tend to compound over time; constant supervision mitigates this problem. Hidden channels are also still allowed to change freely, preventing the first 10 channels from having to simultaneously handle both information propagation and loss minimization, which could otherwise lead to unstable or entangled behavior. However, such training still makes convergence greedy, and the approach could miss solutions that require more complex dynamics.
@@
The models were trained for 800 epochs using the AdamW (Loshchilov and Hutter,, 2019) optimizer, with a learning rate linearly decayed from 0.002 to 0.0001. They have just over 10,000 parameters. In practice, much fewer than 20 hidden channels were needed for a majority of tasks, allowing models smaller than 2500 parameters to perform similarly in some cases.
@@
Indeed, the aim is not to compete with such systems, but to study a fundamentally different approach to problem-solving through local, self-organizing computation. Therefore, a closer examination of tasks will be more insightful than overall performance, as will be illustrated in the next subsection. Finally, on 95 of the attempted tasks, training achieves a categorical cross entropy loss <0.01<0.01, with only half of the successful tasks in this group. An additional 48 tasks had over 90% of pixels in the end grid correct. While these metrics are not perfect signals of success, they further underscore the value of deeper analysis into the behaviors and limitations of this approach.
@@
Figure 3: Spiral pattern generalized to $\times100$ grid, shown at three stages of the update process (t=0, 50, 110). The model was trained only on the small examples from Figure 2 and tested with asynchronous updates on, yet maintains structural coherence across space and time, indicating strong generalization and stability.
Synthesizing a program to output a spiral given these examples would not be trivial, yet this task is surprisingly simple for cellular automata. The system learned a compact, iterative rule that is precise enough to have an effective degree of size invariance, capable of scaling up to a $\times100$ grid (Figure 3). Encoding the solution as local state transitions demonstrates how complexity of a transformation in a NCA differs fundamentally from that in centralized models. A possible explanation for the success on this task is that it is one with a consistent initial state and relatively few transition states that need to be learned. The rules are applied iteratively and there is little structural variation or noise that lead to overfitting. The only difference between inputs is grid size, which mostly only influences how many transition steps are needed before the full grid is settled. It is not always obvious what structures or operations are more naturally and efficiently represented by the properties of a particular self-organizing system, making them an interesting direction for benchmarks like ARC and broader computational modeling.
@@
As stated previously, many of the successfully solved tasks do not rely on asynchronous updates. However, some tasks demonstrate that the regularization provided by asynchronicity can make a difference.
Figure 4: Task ID: 3aa6fb7a. Input-output pairs. The left two are training pairs, and the right pair is the test task. This task is structurally simple, filling in the corner of $\times2$ squares, but one of four orientations does not appear in the training examples, which is a problem for Neural Cellular Automata.
@@
The choice of number of time steps is another factor with noticeable impact. For benchmarking, 10 steps were used for consistency, though this setting is somewhat arbitrary and not optimal for every task. Some tasks, like the spiral, generalized more reliably to larger grids (e.g., $\times100$) when trained with 30 steps, despite identical training examples. This suggests that certain transformations benefit from longer iterative computation, and that fixed step counts may limit what the model can express.
@@
A key strength of this approach is its computational efficiency. Models trained in these experiments required only minutes per task on a single consumer GPU. This contrasts sharply with ARC approaches using large language models or other deep learning architectures, which typically involve extensive pretraining or many hours of computation.
However, direct comparison with these approaches is not straightforward or particularly meaningful due to fundamental differences in assumptions and constraints. Instead of seeing these approaches as competing ideas, it may be more productive to consider how their insights and techniques might integrate with NCAs and potentially address limitations of the current approach.
@@
  • Architectural improvements: Exploring ideas from mainstream neural network research, such as attention, gating, or adaptive computation, that address similar weaknesses in recurrent models but have not yet been widely explored with self-organizing models. Given their structural similarities to other neural architectures, NCAs could benefit from adapting these techniques in ways that preserve local update dynamics.